OpenStudy (toxicsugar22):

a student conduct an expirement by reacting 71.3 g Nal with an excess of AgNO3. The reaction yields 75.940 Agl. What is the percent yield. AgNO3 (aq)+ Nal(aq)= AgI(s)+NaNO3(aq)

OpenStudy (anonymous):

please apply the procedure with the 52.1 value here then you give me the answer to check

OpenStudy (anonymous):

I think I can help!:)

OpenStudy (anonymous):

i made acorrection to the first one with a) and b) so please check it out a bit

OpenStudy (anonymous):

@jadehoff thank you

OpenStudy (anonymous):

Your welcome, Well you have to do what @telijahmed said to do. :)

OpenStudy (anonymous):

pply the procedure with the 52.1 value

OpenStudy (anonymous):

i am on it

OpenStudy (anonymous):

First step calculate the theoritical mass of AgI and this is the procedure calculate the moles of NaI=71.3/23+127 which is 445.623 *10^-3 moles the ratio of AgI to NaI is 1:1 WHICH MEANS THAT THE MOLES OF AgI would also be 445.623 *10^-3 moles calculate now the theoritical mass of AgI and it is given by the formula MASS=MOLES*MOLAR MASS MASS OF AgI = 445.623 *10^-3 *(127+197) the value is 144.3825 grams now that you have THEORITICAL MASS = 144.3825 ACTUAL MASS=75.940 PERCENT YIELD=(ACTUAL M/THEORITICAL MASS) * 100 =(75.940/144.3825)*100=52.5964019%

OpenStudy (anonymous):

52.5964019%

OpenStudy (anonymous):

double check if i fit in the right values and if so then this is the right answer

OpenStudy (anonymous):

please get back to me to questions about the concept because i am about to go on a long break from the forum

OpenStudy (toxicsugar22):

this is great

OpenStudy (toxicsugar22):

thanks